# Evaluation of the absolute single-photon detection efficiency of HRPPD

### v2

@@ 7 HRPPD Collection Efficiency
The Collection Efficiency follows from the PDE and QE measurements at the same wavelength and at the same physical location in the HRPPD active area as CE = PDE/QE. Substituting the numbers obtained in the last two sections, one can see that the HRPPD CE is 70.3 $\pm$ 1.4%, as measured at a particular active area spot. This number looks reasonable given that most of the primary photoelectrons hitting the interstitial space (see Figure 11) between the pores of the top surface of the first MCP either do not bounce back or do not make it into the pores with an energy sufficient to initiate an avalanche. For the MCPs used in EIC HRPPD #27, the pores occupy 69.4 % of the surface area [8], in a reasonable agreement with our CE estimate.
@@ 8 Conclusions
The photon detection efficiency of one HRPPD from the first batch dedicated for EIC has been measured to be 17.13 $\pm$ 0.11 (stat) $\pm$ 0.29 (sys) % at 398.6 nm, in a photoelectron pulse counting mode using a picosecond diode laser. HRPPD QE at the same spot and at the same wavelength has been evaluated to be 24.37 $\pm$ 0.03 (stat) $\pm$ 0.32 (sys) %, leading to a CE estimate of 70.3 $\pm$ 1.6%, which is consistent with the expectations.
@@ Acknowledgements
We are grateful to Dr. Zhoudunming Tu and the Center for Functional Nanomaterials at BNL for their assistance with the scanning electron microscope imaging and to Bob Azmoun for his contribution to the setup commissioning.

### v3

@@ 7 HRPPD Collection Efficiency
The Collection Efficiency follows from the PDE and QE measurements at the same wavelength and at the same physical location in the HRPPD active area as CE = PDE/QE. Substituting the numbers obtained in the last two sections, one can see that the HRPPD CE is (70.3 $\pm$ 1.6)%, as measured at a particular active area spot. This value is reasonable given that most of the primary photoelectrons striking the interstitial space between the pores of the top surface of the first MCP (Figure 11) either do not bounce back or do not make it into the pores with an energy sufficient to initiate an avalanche. For the MCPs used in EIC HRPPD #27, its open area ratio of the entry surface is 69.4% [16], in a reasonable agreement with our CE estimate. Images from scanning electron microscopy reveal that the open area ratio across the MCP entry surface varies by at most 4% for a beam spot with 1 mm diameter.
@@ 8 Conclusions
The photon detection efficiency of one HRPPD from the first batch dedicated for EIC has been measured to be (17.1 $\pm$ 0.1 [stat] $\pm$ 0.3 [sys])% at 398.6 nm, for a pixel near the center, in a photoelectron pulse counting mode using a picosecond diode laser. HRPPD QE at the same spot and at the same wavelength has been evaluated to be (24.4 $\pm$ 0.1 [stat] $\pm$ 0.3 [sys])%, leading to a CE estimate of (70.3 $\pm$ 1.6)%, which is consistent with the expectations.
@@ Acknowledgements
We are grateful to Dr. Zhoudunming Tu and the Center for Functional Nanomaterials at BNL for their assistance with the scanning electron microscope imaging and to Bob Azmoun for his contribution to the setup commissioning. This work was supported in part by the U.S. Department of Energy under Prime Contract No. DE-SC0012704.